OpenStudy (anonymous):

if x and y satisfy the equations xy=1 and 1/y - 3/x = 1/2 then the value of x+y=

Miracrown (miracrown):

Start xy = 1 1/y - 3/x = 1/2 For the latter equation, let us clear the fractions by multiplying by 2xy this gives - 2x - 6y = xy Now let us substitute y = 1/x from the first given equation into this one. 2x - 6*(1/x) = x * (1/x) 2x - 6/x = 1. Multiply through by x. 2x^2 - x - 6 = 0. After bringing the x to the LHS. We can factor this (or use quadratic formula). (2x + 3)(x - 2) = 0 works. Therefore x = 2, and x = -3/2. are solutions Plugging this into y = 1/x gives us two solution coordinates (2,1/2) and (-3/2, -2/3) namely. For the first pair x+y is 5/2. For the second pair x+y is -13/6.

OpenStudy (anonymous):

The option are given are A. -7/8 OR 5/6 B. -2/3 OR -3/4 C.-1/1/4 OR 1/2/3 D. -3/1/3 OR 1/1/2 E. -2/1/6 OR 2/1/2 @Miracrown

Miracrown (miracrown):

Yes, okay, so we have the right answers I am just trying to match that with the representation in the answers you gave E presents them as mixed fractions Know what I mean? @HemarubenY

OpenStudy (anonymous):

sorry ! i didn't notice that .. thank you for ur guide ! @Miracrown

Miracrown (miracrown):

no worries, but do you understand it all? Any doubts? - don't hesitate to share it with me... ;)

OpenStudy (anonymous):

No doubts .. thanks :) @Miracrown